Neil will pass his comment concerning drupal to Craig.
Graham commented that this should read "Find out about access to ...."
After meeting of 10/7, it was decided that this was no longer necessary in the short term and can await account lifecycle code.
Graham to file an enhancement request.
Agreed that Alastair should be part of this discussion. Alison to create wiki page and include comments received so far.
Follow-up required, details of which will be discussed by Craig and Neil. Final report, to include a list of people and responses, should be sent to Dave Robertson.
All Units to update cookie list to show action has been completed.
Craig to update blog article with link to document.
Blog articles suggested:
Blog articles completed:
Report from Computing Executive Group
Reports from units.
Please note that it has now been agreed that when kit needs to be moved between Forum and KB, we should contact the technicians via an RT ticket in the techs queue. They will then arrange the hiring of the van etc. When arranging this, we should check whether anyone else also has kit that they need moved to avoid unnecessary workload for the technicians.
Stephen asked that everyone reboot their machines as soon as possible.
Please let MPU know as soon as possible of any suitable/unsuitable times for the reboot.
Please continue efforts to stop using VMware.
Craig confirmed that queue names will stay the same. We will have to ensure that staff are fully aware of the changes planned for 24th July. We'll need to ensure that MDP users have the correct ppd file. As a result of these changes, scanning will be to email only. There will no longer be a samba option. Subsequent to the report being written, Craig confirmed that the issue with access to ldap had been resolved.
George asked about the situation with upgrading the AFS DB/KDC slave servers. As the AFS component works on sl6, there is no reason why upgrades can't be scheduled. Whether it was appropriate to have the AFS DB servers on the KDCs was discussed and it was agreed that this was OK. We currently have a DB server at each site but this wouldn't be possible with virtual machines. Currently it takes about 2 minutes for a client using a DB server which has disappeared to switch to an alternative server. Stephen mentioned that this could be reduced.
It was agreed, as noted above, to postpone discussing the safe disposal of disks until the next meeting.
The next meeting will be on Wednesday 25th July 2012 in 4.31 chaired by Alison Downie.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh